OpenStudy (anonymous):

@Hero i have another problem just like the last one i think i did something wrong

OpenStudy (anonymous):

FV=PV(1+rt) 4,298.55=1,987.22(1+15r) divide both sides by 1,987.22=2.1630 subtract 1=1.1630 divide by 15 that=0.077 7.7% not even close to the answer im supposed to get

OpenStudy (anonymous):

@ranga can you help me please

OpenStudy (anonymous):

i think im supposed to use compound interest to solve

OpenStudy (anonymous):

yes i think unless i did it wrong. i got 7.7% when my options are 5.2% 0.4% 13.3% 5.8%

OpenStudy (anonymous):

yes the problem says angela currently has an account balance of $4,298.55. She opened the account 15 years ago with a deposit of $1,987.22. If the interest compounds monthly, what is the interest rate on the account?

OpenStudy (anonymous):

i got 21.06

hero (hero):

@Shorthell, you are not paying close enough attention to what equations to use.

hero (hero):

For simple interest, you use the simple interest formula. For compound interest, you use compound interest.

OpenStudy (anonymous):

yes i have tried the compound interest formula. the formula is A=P(1+r/n)^nt

OpenStudy (anonymous):

180 months in 15 years

hero (hero):

@shorthell, you need to use the appropriate formulas for the problem.

OpenStudy (anonymous):

the compound formula correct.

hero (hero):

You are working on Future Value and Present Value. I gave you the formulas earlier. Yes, but it is appropriate to use the formulas in proper context.

OpenStudy (anonymous):

i have went back to that question earlier and tried doing this problem i just cant get it

hero (hero):

All you need is the formulas for FV and PV as they apply to either simple interest or compound interest...both of which I gave you in the last question.

hero (hero):

It's not that difficult. You just apply the proper formula. You won't be able to use just one formula for every problem.

OpenStudy (anonymous):

so i use the FV=PV(1+rt)? or the

hero (hero):

That's for simple interest as I already explained in the last question you posted.

OpenStudy (anonymous):

this is the only other formula we talked about in the last question. fail i mean FV=PV(1+i)^n

hero (hero):

And that is for compound interest. Now which one do you think you are supposed to use here? Re-read the problem again.

OpenStudy (anonymous):

compound

OpenStudy (anonymous):

4,298.5=1,987.22(1+i)^15

hero (hero):

Now, what do you think the t stands for? Years or number of times compounded?

OpenStudy (anonymous):

number of times so instead of 15 would it be 12 or 180

hero (hero):

Number of months compounded in this this case. Which is (12)(15) = 180

OpenStudy (anonymous):

so4,298.5=1,987.22(1+i)^180

hero (hero):

so just to be clear, t = 180

hero (hero):

Yes, correct.

hero (hero):

You know how to solve the rest (I hope)

OpenStudy (anonymous):

divide both sides by 1,987.22

OpenStudy (anonymous):

subtract 1 then divide by 180

hero (hero):

Now, hang on there.

hero (hero):

You have to pay attention to what you are doing.

hero (hero):

I know what we did last time, but look at this set up we have here. The 180 is an exponent, correct?

OpenStudy (anonymous):

correct i see

hero (hero):

So that means we can't just divide by 180

hero (hero):

If you knew logs, your life would be a lot easier

hero (hero):

The truth is, the 1 + r = b

hero (hero):

4,298.5=1,987.22(b)^180

OpenStudy (anonymous):

i am confused. you do divide first correct

hero (hero):

You're confused because you don't know what to do with the exponent.

hero (hero):

There's two ways to solve it. You can either take logs or you do can take the 180th root of both sides.

hero (hero):

You have to learn one of those methods.

OpenStudy (anonymous):

can you teach me how to log?

hero (hero):